# Interacting Multiple Model-based Joint Homography Matrix and Multiple Object State Estimation

### v1

@@ 4.1 State Definition and Motion Models
$\mathbf{x}=\begin{bmatrix}x^{G,x}\\ \dot{x}^{G,x}\\ x^{G,y}\\ \dot{x}^{G,y}\\ \mathbf{H}^{G,1}\\ \mathbf{H}^{G,2}\\ \mathbf{H}^{G,3}\end{bmatrix},$ (11)
where $\mathbf{H}^{G,1}$, $\mathbf{H}^{G,2}$ and $\mathbf{H}^{G,3}$ are the first, second and third columns of the homography matrix, respectively.
Note that although each track is treated as having its own unique projection matrix, in reality, there is a single projection matrix that relates the image plane to the ground plane. Examining Figure 3, this would imply that all of the current tracks’ ground plane positions become dependent upon receiving an image plane measurement. However, this work handles each track independently to simplify track management.
The proposed method makes use of the constant velocity model for the ground plane dimensions as in (5). For the homography state elements, both the dynamics models of (8) and (9) are combined with an IMM filter. The reader is referred to [49, 50, 51] for details regarding IMM filters. The process noise covariance matrix corresponding to the homography state elements is determined as Section 4.3 explains.
@@
$n$ denotes the size of the buffer, which stores the bounding box measurements $\mathbf{x}^{M}_{t-n:t-1}$ previously associated with the particular track. The superscript $M$ indicates a measurement, except in $\mathbf{x}^{M}_{t\mid t-1}$ which refers to the predicted measurement at time $t$ given the measurements from $t-n$ to $t-1$. Each measurement is stored as $\begin{bmatrix}x^{I,l}&x^{I,t}&x^{I,r}&x^{I,b}\end{bmatrix}$, where the components denote the left x-, top y-, right x- and bottom y-coordinate of the corresponding bounding box, respectively. The maximum size of the bounding box measurement buffer is fixed to $n=5$. When fewer than two measurements are in the buffer, the last-known bounding box is used as the prediction.
@@
Figure 4: Showing how coupling bounding box predictions to ground plane state estimates may be beneficial for re-identification. Green bounding boxes represent confirmed tracks, while red represents coasted tracks. The track ID is shown at the top-left above each bounding box. Blue dots represent the ground plane position of the corresponding confirmed track as projected into the image plane, while a red dot represents the same for a coasted track. By coupling the bounding box prediction for track 3 to its predicted ground plane position, it can be re-identified with the BIoU score. The predicted bounding box goes off-screen without this coupling, and the track cannot be re-identified.
@@ 4.2 Measurement Model
Let $\bar{x}^{I,x}$ and $\bar{x}^{I,y}$ denote the bottom-centre x- and y- bounding box coordinates in the image plane as obtained in (6). A track’s ground plane coordinates $x^{G,x}$ and $x^{G,y}$ are related to these image coordinates through its homographic projection state elements:
$\begin{bmatrix}\bar{x}^{I,x}_{t}\\ \bar{x}^{I,y}_{t}\\ 1\end{bmatrix}=\operatorname{norm}\left(\mathbf{H}^{G}_{t}\begin{bmatrix}x^{G,x}_{t}\\ x^{G,y}_{t}\\ 1\end{bmatrix}\right)+\mathbf{R}_{t}=h(\mathbf{x})+\mathbf{R}_{t}.$ (13)
@@
$\mathbf{R}_{0}=\left[\begin{array}[]{cc}\left(\sigma_{m}x^{I,w}_{0}\right)^{2}&0\\ 0&\left(\sigma_{m}x^{I,h}_{0}\right)^{2}\end{array}\right]$ (14)
@@
$\mathbf{H}^{G}_{t}=\begin{bmatrix}h_{1}&h_{2}&h_{3}\\ h_{4}&h_{5}&h_{6}\\ h_{7}&h_{8}&h_{9}\end{bmatrix}.$
Following a similar derivation to that presented in [14], the Jacobian matrix of (13) with respect to the ground plane coordinates is
$\frac{\partial\left(x^{I,x_{g}}_{t},x^{I,y_{g}}_{t}\right)}{\partial\left(x^{G,x}_{t},x^{G,y}_{t}\right)}=\gamma\begin{bmatrix}h_{1}-\frac{b_{1}}{b_{3}}h_{7}&h_{2}-\frac{b_{1}}{b_{3}}h_{8}\\ h_{4}-\frac{b_{2}}{b_{3}}h_{7}&h_{5}-\frac{b_{2}}{b_{3}}h_{8}\end{bmatrix},$ (15)
@@
$\begin{split}&\frac{\partial\left(x^{I,x_{g}}_{t},x^{I,y_{g}}_{t}\right)}{\partial\left(h_{1},h_{4},h_{7},h_{2},h_{5},h_{8},h_{3},h_{6}\right)}=\\ &\gamma\begin{bmatrix}x^{G,x}_{t}&0&-\frac{b_{1}}{b_{3}}x^{G,x}_{t}&x^{G,y}_{t}&0&-\frac{b_{1}}{b_{3}}x^{G,y}_{t}&1&0\\ 0&x^{G,x}_{t}&-\frac{b_{2}}{b_{3}}x^{G,x}_{t}&0&x^{G,y}_{t}&-\frac{b_{2}}{b_{3}}x^{G,x}_{t}&0&1\end{bmatrix}.\end{split}$ (16)
Note that in practice, $\mathbf{H}^{G}_{t}$ is normalised with respect to $h_{9}$ such that $h_{9}=1$. Therefore, the partial derivative with respect to $h_{9}$ is 00.
@@ 4.3 Dynamic Noise Estimation
$\mathbf{Q}_{t}=\mathrm{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{T}\right].$
@@
$\mathbf{R}_{t}=\mathrm{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{T}\right].$
@@
$\mathrm{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{T}\right]=\mathbf{\varepsilon}_{t}\mathbf{\varepsilon}_{t}^{T}+\mathbf{J}_{t}\mathbf{P}^{-}_{t}\mathbf{J}_{t}^{T},$ (17)
where $\mathbf{J}_{t}$ is the Jacobian matrix with respect to the predicted state components obtained from (15), (16). $\mathbf{P}^{-}_{t}$ is the a priori state covariance matrix. Let $\mathbf{\bar{x}}^{M}_{t}$ denote the measured bottom-centre x- and y- bounding box coordinate vector, then $\mathbf{\varepsilon}_{t}=\left[\mathbf{\bar{x}}^{M}_{t}-h\left(\mathbf{x}^{+}_{t}\right)\right]$ is the residual between the measurement and the updated state vector projected into the image plane. This paper makes use of a moving window to average noise covariance estimates such that
@@
where $m=5$. Note that the Kalman filter update step requires an estimate of $\mathbf{R}_{t}$, yet $\mathbf{R}_{t}$ depends on the residual of the updated state vector in (17). To make use of as much information as possible at the current time step, a “dummy” Kalman filter update is first performed to obtain the quantity in (17). The actual update step is then performed with $\mathbf{R}_{t}$ obtained from (18).
@@
where $\mathbf{d}_{t}=\left[\mathbf{\bar{x}}^{M}_{t}-h\left(\mathbf{x}^{-}_{t}\right)\right]$ is the measurement innovation and $\mathbf{x}^{-}_{t}$ is the predicted a priori state estimate. Similar to (18), a moving window is used to average process noise covariance estimates obtained from (19). The process noise covariance matrix for the homography state elements, which will be used in the subsequent prediction at time step $t+1$, is obtained by
$\mathbf{Q}^{G,H}_{t}=\frac{1}{m}\sum_{i=t-m+1}^{t}\mathrm{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{T}\right].$
Only the matrix elements which correspond to homography state elements are extracted from (19), such that the final process covariance matrix for the entire state is $\mathbf{Q}_{t}=\operatorname{diag}(\mathbf{Q}^{G,x},\mathbf{Q}^{G,H}_{t})$. $\mathbf{Q}^{G,H}_{t}$ refers either to $\mathbf{\tilde{Q}}^{G,H}_{t}$ or to $\mathbf{\bar{Q}}^{G,H}_{t}$, depending on the filter motion model in question.
@@ 4.4 Association and Track Management
This section describes the association algorithm and how the BIoU and Mahalanobis distances are combined to form robust association scores between tracks and candidate detections. The Hungarian algorithm (specifically, the Jonker-Volgenant variation [53]) uses these scores in cascaded matching stages to obtain track-detection associations. There are three association stages in total, with association score thresholds $\alpha_{1},\alpha_{2}$ and $\alpha_{3}$, respectively. The values of the hyperparameters introduced in this section are obtained as described in Section 5.
@@ 4.4.1 Track Management
Before the association algorithm is described, it is necessary to define the track states that determine by which branch of the algorithm a specific track will be considered. When a track is initialised, it is in the tentative state. From the tentative state, a track becomes confirmed if it is associated with a detection for two consecutive time steps after initialisation; otherwise, it is deleted if it is not associated with any detection for two consecutive time steps. If a confirmed track’s position is predicted without a measurement update (i.e. it is not associated with any detection), it is coasting and is in the coasted state. If a measurement is associated with a coasted track, its state changes again to confirmed. If a coasted track is not associated with a measurement within $\Omega$ time steps, it is deleted.
@@ 4.4.2 The Association Algorithm
$D=\mathbf{d}_{t}^{T}\mathbf{S}_{t}^{-1}\mathbf{d}_{t}+\ln|\mathbf{S}_{t}|,$
where $\mathbf{S}_{t}=\mathbf{J}_{t}\mathbf{P}_{t}^{-}\mathbf{J}_{t}^{T}+\mathbf{R}_{t-1}$, $|\mathbf{S}_{t}|$ denotes the determinant of $\mathbf{S}_{t}$, $\ln$ is the natural logarithm and $\mathbf{R}_{t-1}$ is the measurement noise covariance estimate (18) obtained in the previous update step. For reasons that will be explained subsequently, $D$ is converted to a probability by
$P(D)=1-\operatorname{CDF}(D),$
where $\operatorname{CDF}$ is the cumulative distribution function of the chi-squared distribution, which is defined to have $24$ degrees of freedom (determined experimentally). As a result, $P(D)$ is in the range $[0,1]$. Finally, $\text{BIoU}(\mathbf{x}^{M}_{t},\mathbf{x}^{M}_{t|t-1})$ denotes the BIoU between a given detection $\mathbf{x}^{M}_{t}$ and the bounding box predicted by (12). The BIoU is calculated as in [12], where the measured and predicted bounding box widths and heights are each scaled by $12b+$ 1; in this paper, only a single buffer scale parameter is used.
@@
$\left(\mu^{I}_{t|t-1}\text{BIoU}(\mathbf{x}^{M}_{t},\mathbf{x}^{M}_{t|t-1})+\mu_{t|t-1}^{G}P(D)\right)d_{\text{conf}},$ (21)
@@
where $\mu^{I}_{t|t-1}$ and $\mu^{G}_{t|t-1}$ refer to the predicted model probabilities of the image and ground plane filters, respectively. These weighting factors are determined by (10). From (21), it should be clear that $D$ is converted to probability $P(D)$ to allow the Mahalanobis distance to be mixed with the BIoU on the same scale.
@@
This association stage manifests the filter architecture depicted in Figure 5. While the ground plane filters are mixed explicitly with an IMM filter in the manner established in the literature [49, 50, 51], the image and ground plane filters are handled independently. Transition probabilities between these filters are only defined to use the predicted model probability (10) during association. Each model probability is initialised as $\mu^{I}_{0}=\mu^{G}_{0}=0.5$. When a measurement is associated with a track, these model probabilities are updated [49]:
$\mu^{i}_{t|t}=\frac{\mu^{i}_{t|t-1}\Lambda^{i}_{t}}{\sum_{j}\mu^{j}_{t|t-1}\Lambda^{j}_{t}},$
@@
where $\Lambda^{i}_{t}$ is the likelihood of the associated measurement given model $i\in\{I,G\}$. Specifically, the proposed method uses $\Lambda^{I}_{t}=\text{BIoU}(\mathbf{x}^{M}_{t},\mathbf{x}^{M}_{t|t-1})$ and $\Lambda^{G}_{t}=P(D)$. This is another reason for converting $D$ to a probability: so that the likelihoods $\Lambda^{I}_{t}$ and $\Lambda^{G}_{t}$ are on the same scale.
@@
The second association stage helps to avoid tracks from coasting when their detections in the image plane diverge from their expected value as predicted by the ground plane filter. This may occur due to partial occlusion or irregular, rapid object motion. It may also occur when ground plane filter state estimates are inaccurate, notably when a track is re-identified after a period during which it was coasted [11]. In these cases, the predicted model probability $\mu^{G}_{t|t-1}$ is expected to be close to zero, while the predicted model probability $\mu^{I}_{t|t-1}$ is expected to be close to one. Thus, the association score (21) will be dominated by the BIoU and the track may still be associated despite low $P(D)$. Note that [11] introduced a re-update step to mitigate the effect of state estimation noise and that the second association stage proposed here could be considered an alternative to their solution.
@@
After the second association stage, the remaining unassociated tracks are placed in the coasted state. Of the unassociated detections, those with detection confidence greater than $d_{\text{high}}$ are used in the third association stage to be associated with the tentative tracks. The third association stage also makes use of the association score in (21). The unassociated detections that remain after this stage are used to initialise new tracks.
@@
High ConfidenceDetectionsConfirmed TracksCoasted TracksStart Time Step $t$ Get Detections Predict All Track Positions Association Stage 1 w. Equation (20) Association Stage 2 w. Equation (21) Low ConfidenceDetectionsRemaining DetectionsRemaining Tracks Measurement Updates Confirmed TracksConfirmed TracksRemaining TracksRemaining DetectionsTrack Older Than $\Omega$?High Confidence Detection?YesDeleteNoMark CoastedNoDeleteTentative Tracks Association Stage 3 w. Equation (21) YesUnassociated for 2 Time Steps? Initiate New Tracks Associated for 2 Time Steps?YesDeleteYesMark ConfirmedMark TentativeRemaining TracksRemaining DetectionsMatched Tracks

### v2

@@ 4.1 State Definition and Motion Models
$\mathbf{x}_{t}=\begin{bmatrix}x^{W}_{t}\\ \dot{x}^{W}_{t}\\ y^{W}_{t}\\ \dot{y}^{W}_{t}\\ \mathbf{h}^{W}_{1}\\ \mathbf{h}^{W}_{2}\\ \mathbf{h}^{W}_{3}\end{bmatrix},$ (13)
where $\mathbf{h}^{W}_{1}$, $\mathbf{h}^{W}_{2}$ and $\mathbf{h}^{W}_{3}$ are the first, second and third columns of $\mathbf{H}^{W}_{t}$, respectively.
Note that although each track is treated as having its own unique projection matrix, in reality, there is a single projection matrix that relates the image plane to the ground plane. Examining Figure 3, this would imply that the current tracks’ ground plane positions become dependent upon receiving an image plane measurement. However, this work handles each track independently to simplify track management.
The proposed method uses the constant velocity model for the ground plane dimensions as in (5). For the homography state elements, both the dynamics models of (10) and (11) are combined with an IMM filter. The reader is referred to [77, 78, 79, 80] for details regarding IMM filters, but an overview is provided in A. The process noise covariance matrix corresponding to the homography state elements is determined as Section 4.3 explains.
@@
$n$ denotes the size of the buffer, which stores the bounding box measurements $\mathbf{x}^{M}_{t-n:t-1}$ previously associated with the particular track. The superscript $M$ indicates a measurement, except in $\mathbf{x}^{M}_{t\mid t-1}$ which refers to the predicted measurement at time $t$ given the measurements from $t-n$ to $t-1$. Each measurement is stored as $\begin{bmatrix}l^{I}&t^{I}&r^{I}&b^{I}\end{bmatrix}$, where the components denote the left x-, top y-, right x- and bottom y-coordinate of the corresponding bounding box, respectively. The maximum size of the bounding box measurement buffer is fixed to $n=5$. When fewer than two measurements are in the buffer, the last-known bounding box is used as the prediction.
@@
Figure 4: Showing how coupling bounding box predictions to ground plane state estimates may benefit re-identification. Green bounding boxes represent confirmed tracks, while red represents coasted tracks. The track ID is shown at the top-left above each bounding box. Blue dots represent the ground plane position of the corresponding confirmed track as projected into the image plane, while a red dot represents the same for a coasted track. By coupling the bounding box prediction for track 3 to its predicted ground plane position, it can be re-identified with the BIoU score. The predicted bounding box goes off-screen without this coupling, and the track cannot be re-identified.
@@ 4.2 Measurement Model
Let $\bar{x}^{I}$ and $\bar{y}^{I}$ denote the bottom-centre x- and y- bounding box coordinates in the image plane as obtained in (8). A track’s ground plane coordinates $x^{W}$ and $y^{W}$ are related to these image coordinates through its homographic projection state elements:
$\begin{bmatrix}\bar{x}^{I}_{t}\\ \bar{y}^{I}_{t}\\ 1\end{bmatrix}=\operatorname{norm}\left(\mathbf{H}^{W}_{t}\begin{bmatrix}x^{W}_{t}\\ y^{W}_{t}\\ 1\end{bmatrix}\right)+\mathbf{R}_{t}=h(\mathbf{x})+\mathbf{R}_{t},$ (15)
@@
$\mathbf{R}_{0}=\left[\begin{array}[]{cc}\left(\sigma_{m}w^{I}_{0}\right)^{2}&0\\ 0&\left(\sigma_{m}h^{I}_{0}\right)^{2}\end{array}\right]$ (16)
@@
$\mathbf{H}^{W}_{t}=\begin{bmatrix}h_{1}&h_{2}&h_{3}\\ h_{4}&h_{5}&h_{6}\\ h_{7}&h_{8}&h_{9}\end{bmatrix}.$ (18)
As shown in B, the Jacobian matrix of (15) with respect to the ground plane coordinates is
$\frac{\partial\left(x^{I}_{t},y^{I}_{t}\right)}{\partial\left(x^{W}_{t},y^{W}_{t}\right)}=\gamma\begin{bmatrix}(h_{1}-h_{7}x^{I}_{t})&(h_{2}-h_{8}x^{I}_{t})\\ (h_{4}-h_{7}y^{I}_{t})&(h_{5}-h_{8}y^{I}_{t})\end{bmatrix},$ (19)
@@
$\begin{split}&\frac{\partial\left(x^{I}_{t},y^{I}_{t}\right)}{\partial\left(h_{1},h_{4},h_{7},h_{2},h_{5},h_{8},h_{3},h_{6}\right)}=\\ &\gamma\begin{bmatrix}x^{W}_{t}&0&-x^{I}_{t}x^{W}_{t}&y^{W}_{t}&0&-x^{I}_{t}y^{W}_{t}&1&0\\ 0&x^{W}_{t}&-y^{I}_{t}x^{W}_{t}&0&y^{W}_{t}&-y^{I}_{t}y^{W}_{t}&0&1\end{bmatrix}.\end{split}$ (20)
In practice, $\mathbf{H}^{W}_{t}$ is normalised with respect to $h_{9}$ such that $h_{9}=1$. Therefore, the partial derivative with respect to $h_{9}$ is 00.
@@ 4.3 Dynamic Noise Estimation
$\mathbf{Q}_{t}=\mathrm{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{T}\right].$ (21)
@@
$\mathbf{R}_{t}=\mathrm{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{T}\right].$ (22)
@@
$\mathrm{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{T}\right]=\mathbf{\varepsilon}_{t}\mathbf{\varepsilon}_{t}^{T}+\mathbf{J}_{t}\mathbf{P}_{t|t-1}\mathbf{J}_{t}^{T},$ (23)
where $\mathbf{J}_{t}$ is the Jacobian matrix with respect to the predicted state components obtained from (19), (20). $\mathbf{P}_{t|t-1}$ is the a priori state covariance matrix. Let $\mathbf{\bar{x}}^{M}_{t}$ denote the measured bottom-centre x- and y- bounding box coordinate vector, then $\mathbf{\varepsilon}_{t}=\left[\mathbf{\bar{x}}^{M}_{t}-h\left(\mathbf{\hat{x}}_{t|t}\right)\right]$ is the residual between the measurement and the updated state vector estimate $\mathbf{\hat{x}}_{t|t}$ projected into the image plane. This paper makes use of a moving window to average noise covariance estimates such that
@@
where $m=5$. Note that the Kalman filter update step requires an estimate of $\mathbf{R}_{t}$, yet $\mathbf{R}_{t}$ depends on the residual of the updated state vector in (23). To use as much information as possible at the current time step, a “dummy” Kalman filter update is first performed to obtain the quantity in (23). The actual update step is then performed with $\mathbf{R}_{t}$ obtained from (24).
@@
where $\mathbf{d}_{t}=\left[\mathbf{\bar{x}}^{M}_{t}-h\left(\mathbf{\hat{x}}_{t|t-1}\right)\right]$ is the measurement innovation and $\mathbf{\hat{x}}_{t|t-1}$ is the predicted a priori state estimate. Similar to (24), a moving window is used to average process noise covariance estimates obtained from (25). The process noise covariance matrix for the homography state elements, which will be used in the subsequent prediction at time step $t+1$, is obtained by
$\mathbf{Q}^{W,\mathbf{H}}_{t}=\frac{1}{m}\sum_{i=t-m+1}^{t}\mathrm{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{T}\right].$ (26)
Only the matrix elements which correspond to homography state elements are extracted from (25), such that the final process covariance matrix for the entire state is $\mathbf{Q}_{t}=\operatorname{diag}(\mathbf{Q}^{W,\mathbf{x}},\mathbf{Q}^{W,\mathbf{H}}_{t})$. $\mathbf{Q}^{W,\mathbf{H}}_{t}$ refers either to $\mathbf{\tilde{Q}}^{W,\mathbf{H}}_{t}$ or to $\mathbf{\bar{Q}}^{W,\mathbf{H}}_{t}$, depending on the filter motion model in question.
@@ 4.4 Association and Track Management
This section describes the association algorithm and how the BIoU and Mahalanobis distances are combined to form robust association scores between tracks and candidate detections. The Hungarian algorithm (specifically, the Jonker-Volgenant variation [82]) uses these scores in cascaded matching stages to obtain track-detection associations. There are three association stages, with association score thresholds $\alpha_{1},\alpha_{2}$ and $\alpha_{3}$, respectively. The values of the hyperparameters introduced in this section are obtained as described in Section 5.
@@ 4.4.1 Track Management
Before the association algorithm is described, it is necessary to define the track states that determine by which branch of the algorithm a specific track will be considered. When a track is initialised, it is in the tentative state. From the tentative state, a track becomes confirmed if associated with a detection for two consecutive time steps after initialisation; otherwise, it is deleted if it is not associated with any detection for two consecutive time steps. If a confirmed track’s position is predicted without a measurement update (i.e. it is not associated with any detection), it is coasting and is in the coasted state. If a measurement is associated with a coasted track, its state changes again to confirmed. If a coasted track is not associated with a measurement within $\Omega$ time steps, it is deleted.
@@ 4.4.2 The Association Algorithm
$D=\mathbf{d}_{t}^{T}\mathbf{S}_{t}^{-1}\mathbf{d}_{t}+\ln|\mathbf{S}_{t}|,$ (28)
where $\mathbf{S}_{t}=\mathbf{J}_{t}\mathbf{P}_{t|t-1}\mathbf{J}_{t}^{T}+\mathbf{R}_{t-1}$, $|\mathbf{S}_{t}|$ denotes the determinant of $\mathbf{S}_{t}$, $\ln$ is the natural logarithm and $\mathbf{R}_{t-1}$ is the measurement noise covariance estimate (24) obtained in the previous update step. For reasons that will be explained subsequently, $D$ is converted to a probability by
$P(D)=1-\operatorname{CDF}(D),$ (29)
where $\operatorname{CDF}$ is the cumulative distribution function of the chi-squared distribution, which is defined to have $24$ degrees of freedom (determined experimentally). As a result, $P(D)$ is in the range $[0,1]$. Finally, $\text{BIoU}(\mathbf{x}^{M}_{t},\mathbf{x}^{M}_{t|t-1})$ denotes the BIoU between a given detection $\mathbf{x}^{M}_{t}$ and the bounding box predicted by (14). The BIoU is calculated as in [12], where the IoU is calculated after the measured and predicted bounding box widths and heights are each scaled by $12b+$ 1; in this paper, only a single buffer scale parameter, $b$, is used.
@@
$\left(\mu^{I}_{t|t-1}\text{BIoU}(\mathbf{x}^{M}_{t},\mathbf{x}^{M}_{t|t-1})+\mu_{t|t-1}^{W}P(D)\right)d_{\text{conf}},$ (30)
@@
where $\mu^{I}_{t|t-1}$ and $\mu^{W}_{t|t-1}$ refer to the predicted model probabilities of the image and ground plane filters, respectively. These weighting factors are determined by (12). From (30), it should be clear that $D$ is converted to probability $P(D)$ to allow the Mahalanobis distance to be mixed with the BIoU on the same scale.
@@
This association stage manifests the filter architecture depicted in Figure 5. While the ground plane filters are mixed explicitly with an IMM filter as established in the literature [80, 77, 78, 79] and explained in A, the image and ground plane filters are handled independently. Transition probabilities between these filters are only defined to use the predicted model likelihood (12) during association. Each model probability is initialised as $\mu^{I}_{0}=\mu^{W}_{0}=0.5$. When a measurement is associated with a track, these model probabilities are updated [77]:
$\mu^{i}_{t}=\frac{\mu^{i}_{t|t-1}\Lambda^{i}_{t}}{\sum_{j}\mu^{j}_{t|t-1}\Lambda^{j}_{t}},$ (31)
@@
where $\Lambda^{i}_{t}$ is the likelihood of the associated measurement given model $i\in\{I,W\}$. Specifically, the proposed method uses $\Lambda^{I}_{t}=\text{BIoU}(\mathbf{x}^{M}_{t},\mathbf{x}^{M}_{t|t-1})$ and $\Lambda^{W}_{t}=P(D)$. This is another reason for converting $D$ to a probability: so that the likelihoods $\Lambda^{I}_{t}$ and $\Lambda^{W}_{t}$ are on the same scale.
@@
The second association stage helps to avoid tracks from coasting when their detections in the image plane diverge from their expected value as predicted by the ground plane filter. This may occur due to partial occlusion or irregular, rapid object motion. It may also occur when ground plane filter state estimates are inaccurate, notably when a track is re-identified after a period during which it was coasted [11]. In these cases, the predicted model likelihood $\mu^{W}_{t|t-1}$ is expected to be close to zero, while the predicted model likelihood $\mu^{I}_{t|t-1}$ is expected to be close to one. Thus, the association score (30) will be dominated by the BIoU, and the track may still be associated despite low $P(D)$. Note that [11] introduced a re-update step to mitigate the effect of state estimation noise and that the second association stage proposed here could be considered an alternative to their solution.
The image and ground plane filters are not mixed explicitly because a target’s ground plane state estimate should not be significantly influenced when the BIoU dominates the association score since this often implies motion away from the ground plane (e.g. a jump) or a partial detection, both of which are expected to result in an inaccurate measurement of the ground plane position. However, if the primary goal is to reduce the effect of state estimation noise as in [11], mixing ground and image plane state estimates may be beneficial.
@@
After the second association stage, the remaining unassociated tracks are placed in the coasted state. Of the unassociated detections, those with detection confidence greater than $d_{\text{high}}$ are used in the third association stage to be associated with the tentative tracks. The third association stage also uses the association score in (30). The unassociated detections that remain after this stage are used to initialise new tracks.
@@
High ConfidenceDetectionsConfirmed TracksCoasted TracksStart Time Step $t$ Get Detections Predict All Track Positions Association Stage 1 w. Equation 27 Association Stage 2 w. Equation 30 Low ConfidenceDetectionsRemaining DetectionsRemaining Tracks Measurement Updates Confirmed TracksConfirmed TracksRemaining TracksRemaining DetectionsTrack Older Than $\Omega$?High Confidence Detection?YesDeleteNoMark CoastedNoDeleteTentative Tracks Association Stage 3 w. Equation 30 YesUnassociated for 2 Time Steps? Initiate New Tracks Associated for 2 Time Steps?YesDeleteYesMark ConfirmedMark TentativeRemaining TracksRemaining DetectionsMatched Tracks